OpenStudy (anonymous):

Erin has developed a new lotion that will help children with sensitive skin. She invests $15,225 in equipment to manufacture and package her lotion. Each bottle costs $3.25 to manufacture and sells for $12. How many bottles of lotion must she make and sell before her business breaks even?

OpenStudy (anonymous):

we will take all the numbers given to us, and plug them into the 'line equation' y = m*x + b no mater now much Erin sells, she has a minimum cost of 15,255 dollars, so let c = 15,255 lets have 'x' be the number of bottles sold. and we will have 'm' be the money she makes on selling each bottle. the money made from each bottle will be the price sold for minus the cost to make the bottle. so, m = 12 - 3.25 now we have all the variables, we can put it into the equation y = m*x + b can you do this step for me?